# Unleashing Multi-Hop Reasoning Potential in Large Language Models through Repetition of Misordered Context

### v1

@@ Abstract
Multi-hop reasoning, which requires multi-step reasoning based on the supporting documents within a given context, remains challenging for large language models (LLMs). LLMs often struggle to filter out irrelevant documents within the context, and their performance is sensitive to the position of supporting documents within that context. In this paper, we identify an additional challenge: LLMs’ performance is also sensitive to the order in which the supporting documents are presented. We refer to this as the misordered context problem. To address this issue, we propose a simple yet effective method called context repetition (CoRe), which involves prompting the model by repeatedly presenting the context to ensure the supporting documents are presented in the optimal order for the model. Using CoRe, we improve the F1 score by up to 30%p on multi-hop QA tasks and increase accuracy by up to 70%p on a synthetic task. Additionally, CoRe helps mitigate the well-known “lost-in-the-middle” problem in LLMs and can be effectively combined with retrieval-based approaches utilizing Chain-of-Thought (CoT) reasoning.
@@
Sangwon Yu1        Ik-hwan Kim1        Jongyoon Song1 Saehyung Lee1        Junsung Park1        Sungroh Yoon1,2thanks:   Corresponding author 1Department of Electrical and Computer Engineering, Seoul National University 2Interdisciplinary Program in Artificial Intelligence, Seoul National University {dbtkddnjs96, 12kimih, coms1580, halo8218, jerryray, sryoon}@snu.ac.kr
@@ 1 Introduction
During this process, the model faces several challenges in terms of grounding to the context (Gao et al., 2023). First, the model struggles to filter out noisy documents that are irrelevant to the correct answer (Shi et al., 2023). In fact, noisy documents can significantly degrade the reasoning performance of LLMs (Cuconasu et al., 2024; Wu et al., 2024). Additionally, LLMs are highly sensitive to the position of supporting documents within the context. Specifically, when a document is positioned in the middle of the context, the model may fail to recognize it properly, resulting in a “lost-in-the-middle” problem (Liu et al., 2023). These issues become more critical as the length of the context grows and the number of retrieved documents increases (Lee et al., 2024; Li et al., 2024).
@@
In this study, we address additional context grounding problems related to multi-hop reasoning that have not been deeply discussed so far. In a multi-hop reasoning task, there are multiple supporting documents, and the model derives the answer by considering each document through step-by-step reasoning. Due to the nature of the decoder-only structure, which performs causal attention, LLMs can only reference documents in a left-to-right order. If this order is presented unfavorably for the model’s reasoning, the model’s performance may be significantly degraded. Following this rationale, we observe that, even in ideal situations without noisy documents, LLMs exhibit up to a 26%p difference in F1 score depending on the order of supporting documents. Since there is no guarantee that supporting documents will be presented to the model in the optimal order during actual tasks, this problem inevitably arises. We call it the misordered context problem. Figure 1 shows an example of the misordered context problem.
@@
To address the problem, we approach it from the perspective of context augmentation. Specifically, by repeating the given context multiple times, we construct an augmented context that allows the supporting documents to be presented to the model in the optimal order. We theoretically explain that by repeating the context containing $k$ supporting documents $k$ times, all possible orders of presenting the documents can be covered. That is, the $k$-repeated context can always present the supporting documents to the model in the optimal order. Based on this theoretical analysis, we propose the context repetition (CoRe) method to tackle the misordered context problem.
@@
We conduct experiments on 3 multi-hop QA benchmarks—HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and MuSiQue (Trivedi et al., 2022)—as well as an additional synthetic task. In the synthetic task, the model is required to search for elements in lists composed of non-negative integer elements, based on context where information is presented in a reverse order. We show that CoRe significantly improves multi-hop QA task performance across various LLMs. For example, in the 2WikiMultihopQA task using the Llama-3.1-8B-Instruct model, we observe up to 30%p improvement in F1 score, and in the synthetic task, we observe up to a 70%p improvement in accuracy. Additionally, empirical analysis demonstrates that CoRe provides performance benefits to the model by improving order-related reasoning. Furthermore, CoRe alleviates the previously known “lost-in-the-middle” problem and can be effectively combined with Chain-of-Thought (CoT) methods in retrieval-augmented tasks.
@@
In summary, our work (1) introduces the misordered context problem, a critical challenge for LLMs in multi-hop reasoning, (2) theoretically proposes an approach of context augmentation to tackle this issue, which leads to our simple but practical CoRe method, and (3) demonstrates that this method can significantly improve model performance across various tasks.
@@ 3 CoRe: Context Augmentation for misordered Supporting Documents
In this section, we empirically demonstrate that the performance of LLMs in multi-hop reasoning is sensitive to the order of supporting documents within the context, a phenomenon we refer to as the misordered context problem. For this issue, we theoretically formalize the problem from the perspective of context structure. Building on this formalization, we propose our method, context repetition, which ensures that supporting documents are presented to the model in an optimal order for reasoning.
@@ 3.1 Problem of misordered Context in Multi-Hop Reasoning
Based on these findings, we demonstrate the answer to the research question: the multi-hop reasoning performance of LLMs is indeed significantly sensitive to the order in which supporting documents are presented. This aligns with the rationale that decoder-only LLMs, operating with causal attention, interpret the context in a fixed left-to-right order. In real-world scenarios, common prompting methods such as retrieval or search-based approaches do not guarantee that supporting documents will be presented to the model in an optimal order. Consequently, the model is often exposed to a misordered context that is disadvantageous for reasoning, leading to degraded performance. We term this issue the misordered context problem, and in the following sections, we describe our approach to addressing this problem.
@@ 3.3 Context Repetition
Here, $C$ represents the given context, and $\oplus$ denotes the concatenation operation. Hence, the function $f_{\text{rep}}^{(k)}(C)$ creates an augmented context by just repeating the given context $k$ times. The following theorem provides an important property of $f_{\text{rep}}^{(k)}(\cdot)$ in relation to $\sigma$:
@@
Based on Corollary 1.1, we conclude that $f_{\text{rep}}^{(k)}(\cdot)$ is an appropriate augmentation function that satisfies Eq. 2. We propose to use this function as a method for solving the misordered context problem. Specifically, we introduce the context repetition (CoRe) methodology, which repeatedly presents the given context to help the model recognize the supporting documents in the optimal order. Figure 3 illustrates the detailed operation of our methodology. In practice, we implemented this method by utilizing a chat template to ensure that the model naturally repeats the context given by the user.
@@ 4 Experiments
In this section, we evaluate CoRe on multi-hop QA tasks and a synthetic task that requires complex reasoning. In addition to the main experiments, we conduct an analysis on the mechanism through which CoRe achieves performance improvements in the model.
Models HotpotQA 2WikiMultihopQA MuSiQue
Llama-3.1-8B 49.31 61.19 38.10 55.77 40.75 34.86 32.05 26.24 27.06
+ CoRe 46.93 66.82 48.15 58.13 44.81 65.11 39.10 38.23 34.08
Mistral-7B 34.42 41.14 18.82 32.90 29.77 15.43 15.14 10.31 11.00
+ CoRe 30.41 45.79 27.78 35.67 36.39 29.09 20.6 15.74 12.36
Qwen2.5-7B 61.96 61.66 41.36 65.47 37.26 39.45 33.56 34.87 37.49
+ CoRe 67.03 70.33 54.20 70.30 42.40 64.54 44.76 43.1 45.08
Phi-3.5-mini 26.44 44.17 27.70 45.26 27.00 35.19 17.60 16.80 12.67
+ CoRe 23.29 49.67 37.64 50.92 32.34 36.71 28.60 26.31 26.59
GPT-4o-mini 66.92 68.66 51.96 76.63 43.84 60.96 45.26 42.04 40.04
+ CoRe 64.72 72.67 57.43 79.14 52.03 66.96 52.47 49.47 42.53
@@ 4.1 Experimental Settings
We conduct experiments on 3 multi-hop QA benchmarks: HotpotQA, 2WikiMultihopQA, and MuSiQue. For HotpotQA and 2WikiMultihopQA, reasoning type annotations are provided for each query. For these 2 datasets, we create evaluation sets by sampling an equal number of queries from each reasoning type within the validation set. In the case of MuSiQue, as described in Section 3.1, query types are categorized by the number of hops. We use the entire validation set as the evaluation set for this dataset. To construct the context of multiple documents, we combine the supporting documents with distracting (noisy) documents assigned to each sample in the datasets. Detailed statistics for each evaluation set are provided in Appendix B.
In addition to the multi-hop QA tasks, we also evaluate model performance on an additional synthetic task. Specifically, we provide the model information about a list that contains non-negative integers as elements in a chained form consisting of two consecutive elements. Afterward, we ask the model for the first element of the list that contains a specific element. At this point, the specific element is the last element of a list. For example, when there are two lists with 3 elements each, the context and question are structured as follows:
@@
Figure 4: Main results in the synthetic task.
Since the elements of each list are presented in sequential order from the first to the last, the process of finding the list that contains the last element and then searching for the first element of that list must be performed in a right-to-left order. This poses a challenge for models that typically process context in a left-to-right order. Moreover, since this type of task does not involve specific factual knowledge, it minimizes the interference of the model’s pre-existing factual knowledge (Xiong et al., 2024). Thus, these characteristics of the synthetic task enable us to establish an environment that clearly exposes the model’s inherent misordered context problem. For our main evaluation, we create a synthetic dataset consisting of 1000 samples. Each sample is composed of 10 lists, and each list contains 3 integers between 0 and $9$ that do not overlap with each other.
We conduct experiments using a total of 5 models, including GPT-4o-mini (OpenAI, 2024), in addition to the 4 LLMs used in section 3.1. For Multi-hop QA tasks, we evaluate by measuring the F1 score based on the short answers generated by the models, and for synthetic tasks, we evaluate by measuring accuracy based on whether the answers generated by the models match the correct answers. All prompts used in the experiments can be found in the Appendix C.
@@ 4.2 Main Results
Table 1 shows the F1 scores of baseline models and models with CoRe applied for multi-hop QA tasks. Overall, performance improvements were observed across all models and datasets. In the case of HotpotQA, a significant performance improvement was seen in the bridge type. This is because answering bridge-type questions requires reasoning across documents by linking information between them, which is closely related to the order of documents. For 2WikiMultihopQA, a large performance improvement was observed in the bridge-comparison type, with Llama achieving a 30-point improvement and Qwen achieving a 25-point improvement in F1 score. This is because the bridge-comparison type is a complex 4-hop reasoning task that combines both bridge and comparison from HotpotQA. For the MuSiQue dataset, overall performance improved regardless of the number of hops. Taken together, these results indicate that CoRe contributes significantly to improving the model’s multi-hop reasoning performance, especially in tasks sensitive to the order of supporting documents.
@@
We measure the accuracy of each model as we increase the number of context repetitions up to 10 in the synthetic task. As shown in Figure 4, we can see that the accuracy when no repetition is performed has a low value between 0% and 60%. This demonstrates that existing LLMs are unable to properly perform the synthetic task with simple text inference alone. However, when CoRe is applied, we observe a significant performance improvement, with accuracy increasing by at least 30% points and up to 70% points. Additionally, we can see that many models show a tendency for performance to improve as the number of repetitions increases. This suggests that as repetition progresses multiple times, the model is finding the optimal order. However, for some models, particularly Qwen, there is a tendency for performance to drop again as the repetition steps increase. This appears to be due to noisy information in the context being repeated, negatively affecting its reasoning ability. We provide a detailed analysis of this in Section 4.3.

### v2

@@ Abstract
Multi-hop reasoning, which requires multi-step reasoning based on the supporting documents within a given context, remains challenging for large language models (LLMs). LLMs often struggle to filter out irrelevant documents within the context, and their performance is sensitive to the absolute position of supporting documents within that context. In this paper, we identify an additional challenge: LLMs’ performance is also sensitive to the order, relative position, in which the supporting documents are presented. We refer to this as the misordered context problem. To address this issue, based on the theoretical approach, we propose a simple yet effective method called context repetition (CoRe), which involves prompting the model by repeatedly presenting the context. This ensures that certain contiguous reasoning segments within supporting documents are presented in the optimal order, effectively guiding the model’s reasoning in the appropriate direction. Applying CoRe, we improve the F1 score by up to 30%p on multi-hop QA tasks and increase accuracy by up to 70%p on a synthetic task. Additionally, CoRe helps mitigate the well-known “lost-in-the-middle” problem in LLMs and can be effectively combined with retrieval-based approaches utilizing Chain-of-Thought (CoT) reasoning.
@@
Sangwon Yu1        Ik-hwan Kim1        Jongyoon Song1 Saehyung Lee1        Junsung Park1        Sungroh Yoon1,2thanks:   Corresponding author 1Department of Electrical and Computer Engineering, Seoul National University 2AIIS, ASRI, INMC, ISRC, and IPAI, Seoul National University {dbtkddnjs96, 12kimih, coms1580, halo8218, jerryray, sryoon}@snu.ac.kr
@@ 1 Introduction
During this process, the model faces several challenges in terms of grounding to the context (Gao et al., 2023). First, the model struggles to filter out noisy documents that are irrelevant to the correct answer (Shi et al., 2023). In practice, noisy documents can significantly degrade the reasoning performance of LLMs (Cuconasu et al., 2024; Wu et al., 2024). Additionally, LLMs are highly sensitive to the absolute position of supporting documents within the context. Specifically, when a document is positioned in the middle of the context, the model may fail to recognize it properly, resulting in a “lost-in-the-middle” problem (Liu et al., 2023). These issues become more critical as the maximum length of the context grows and the number of retrieved documents increases (Lee et al., 2024; Li et al., 2024).
@@
In this study, we address additional context grounding problems related to multi-hop reasoning that have not been deeply discussed so far. In a multi-hop reasoning task, there are multiple supporting documents, and the model derives the answer by considering each document through step-by-step reasoning. Due to the nature of the decoder-only structure, which performs causal attention, LLMs can only perceive documents in a left-to-right order. If this order is presented unfavorably for the model’s reasoning, the model’s performance may be significantly degraded. Following this rationale, we observe that, even in ideal situations without noisy documents, LLMs exhibit up to a 26%p difference in F1 score depending on the order, relative position, of supporting documents. Since there is no guarantee that supporting documents will be presented to the model in the appropriate order for the model reasoning during actual tasks, it is inevitable that this problem arises. We call it the misordered context problem. Figure 1 shows an example of the misordered context problem.
@@
To address the problem, we approach it from the perspective of context augmentation. We aim to construct the augmented context that presents the supporting documents in a more appropriate order than the original context. Through theoretical analysis, we demonstrate that by repeating the context containing $k$ supporting documents $k$ times, all possible orders of presenting the documents can be covered. That is, the $k$-repeated context can always present the supporting documents to the model in the optimal order. In real-world scenarios, however, the value of $k$ for a given context is unknown. Furthermore, repeating the context leads to an increase in inference time costs. Based on this theoretical analysis and its identified limitations, we propose the context repetition (CoRe) method: the given context is repeated $\hat{k}$ times in which $\hat{k}$ is the predetermined hyperparameter. This practical approach ensures that all $\hat{k}$-contiguous reasoning chains within supporting documents are presented in the optimal order, guiding LLM’s reasoning in a more coherent and structured manner.
@@
To evaluate CoRe, we conduct experiments on 3 multi-hop QA benchmarks—HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), and MuSiQue (Trivedi et al., 2022)—as well as an additional synthetic task. In the synthetic task, the model is required to search for elements in lists composed of non-negative integer elements, based on context where information is presented in a reverse order. In the evaluation, CoRe significantly improves multi-hop QA task performance across various LLMs. For example, in the 2WikiMultihopQA task using the Llama-3.1-8B-Instruct model, we observe up to 30%p improvement in F1 score, and in the synthetic task, we observe up to a 70%p improvement in accuracy. Through in-depth empirical analysis, we demonstrate that CoRe provides performance benefits to the model by improving order-related reasoning. Additionally, we show that CoRe alleviates the previously known “lost-in-the-middle” problem and can be effectively combined with Chain-of-Thought (CoT) methods in retrieval-augmented tasks.
@@
In summary, our work (1) introduces the misordered context problem, a critical challenge for LLMs in multi-hop reasoning, (2) theoretically proposes context augmentation as a solution to this issue, which forms the foundation of our practical and effective CoRe method, and (3) demonstrates that CoRe significantly enhances model performance across various tasks and effectively mitigates the misordered context problem.
@@ 2 Related Work
From the perspective of context sensitivity in LLMs, recent works focus on addressing the issue of absolute positional bias in tasks where LLMs are required to choose between multiple options (Pezeshkpour and Hruschka, 2024; Zheng et al., 2024). Also, Liu et al. (2023) also address the issue of absolute positional bias, specifically the "lost in the middle" problem. This problem pertains to the model’s difficulty in effectively recognizing information that appears in the middle positions of the context provided to the LLM. This absolute positional bias issue differs from the ordering problem we address in our work, which involves ensuring that LLMs interpret supporting documents within the context in the correct reasoning chain order, relative position, in the multi-hop reasoning task. Chen et al. (2024) analyze the sensitivity of reasoning performance in LLMs to the order in which rules or premises are presented during logical and mathematical reasoning. Our work investigates longer-context scenarios and addresses a broader spectrum of multi-hop reasoning tasks, including open-domain question answering and synthetic tasks, where document-level ordering issues are more pronounced. Furthermore, we propose a concrete methodology to mitigate order-related challenges, enhancing the novelty of our approach in addressing this problem.
@@ 3 CoRe: Context Augmentation for Misordered Supporting Documents
In this section, we empirically demonstrate that the performance of LLMs in multi-hop reasoning is sensitive to the order of supporting documents within the context. We refer to this phenomenon as the misordered context problem. We theoretically formalize this problem from the perspective of context structure. Building on the formalization and considering its practical limitation, we propose context repetition (CoRe). This method enables LLMs to interpret specific contiguous reasoning segments in an optimal order, facilitating more effective multi-hop reasoning within a given context.
@@ 3.1 Problem of misordered Context in Multi-Hop Reasoning
Based on these findings, we demonstrate the answer to the research question: the multi-hop reasoning performance of LLMs is indeed significantly sensitive to the order in which supporting documents are presented. This also aligns with the rationale that decoder-only LLMs, operating with causal attention, interpret the context in a fixed left-to-right order. In real-world scenarios, common prompting methods such as retrieval or search-based approaches do not guarantee that supporting documents will be presented to the model in an optimal order. Consequently, the model is often exposed to a misordered context that is disadvantageous for reasoning, leading to degraded performance. We term this issue the misordered context problem, and in the following sections, we describe our approach to addressing this problem.
@@ 3.3 Context Repetition
Here, $C$ represents the given context, and $\oplus$ denotes the concatenation operation. Hence, the function $f_{\text{rep}}^{(k)}(C)$ creates an augmented context by just repeating the given context $k$ times. Please note that $k=1$ indicates the absence of repetition. The following theorem provides an important property of $f_{\text{rep}}^{(k)}(\cdot)$ in relation to $\sigma$:
@@
Based on Corollary 1.1, we conclude that $f_{\text{rep}}^{(k)}(\cdot)$ is an appropriate augmentation function that satisfies Eq. 2.
3.4 Practical Approach
From the previous section 3.3, we can conclude that $f_{\text{rep}}^{(k)}(\cdot)$ can serve as a methodology for mitigating the misordered context problem. However, for direct application, the following two issues must be addressed:
  1. 1.
    It is necessary to know the number of supporting documents in the given context, $k$. (Issue 1)
  2. 2.
    The model should effectively handle the computational cost induced by the increased length when the context is repeated $k$ times. (Issue 2)
In real-world multi-hop reasoning scenarios, it is not feasible to fully address issue 1. Additionally, issue 2 becomes intractable as the number of documents included in the context grows, and more importantly, it incurs a significant additional cost in terms of memory.
To overcome these critical limitations, we replace $k$ with a predetermined hyperparameter $\hat{k}$, resulting the function $f_{\text{rep}}^{(\hat{k})}(\cdot)$ which can be considered an approximation of $f_{\text{rep}}^{(k)}(\cdot)$. This directly addresses the issue 1 by constraining the unknown value of $k$ to predetermined $\hat{k}$. While this approach does not guarantee that all supporting documents will be recognized in the optimal order, it ensures that at least the $\hat{k}$-contiguous reasoning chains within supporting documents are interpreted in the optimal order. This allows the LLM to perform multi-hop reasoning from the optimal starting point, guiding the model to execute subsequent reasoning steps correctly. From the perspective of the issue 2, the introduction of $\hat{k}$ significantly reduces the additional computational and memory costs, bringing them to a manageable level.
Consequently, we propose to use $f_{\text{rep}}^{(\hat{k})}(\cdot)$ as a method for solving the misordered context problem. In other words, we introduce the context repetition (CoRe) methodology, which repeatedly presents the given context to guide the model recognize the supporting documents in the appropriate order. Figure 3 illustrates the detailed operation of our methodology when $\hat{k}=2$. In practice, we implement this method by utilizing a chat template to ensure that the model naturally repeats the context given by the user.
@@ 4 Experiments
In this section, we evaluate CoRe on multi-hop QA tasks and a synthetic task that requires complex reasoning. In addition to the main experiments, we conduct an analysis of the mechanism through which CoRe achieves performance improvements in the model.
@@ 4.1 Experimental Settings
We conduct experiments on three multi-hop QA benchmarks: HotpotQA, 2WikiMultihopQA, and MuSiQue. For HotpotQA and 2WikiMultihopQA, reasoning-type annotations are provided for each query. For these two datasets, we create evaluation sets by sampling an equal number of queries from each reasoning type within the validation set. In the case of MuSiQue, as described in Section 3.1, query types are categorized by the number of hops. We use the entire validation set as the evaluation set for this dataset. To construct the context of multiple documents, we combine the supporting documents with distracting (noisy) documents assigned to each sample in the datasets. Detailed statistics for each evaluation set are provided in Appendix B. In the multi-hop QA evaluation, we set the $\hat{k}$ within the range of $[1,3]$ ($\hat{k}=1$ means the naive LLM baseline without repetition).
In addition to the multi-hop QA tasks, we evaluate model performance on an additional synthetic task. Specifically, we provide the model information about a list that contains non-negative integers as elements in a chained form consisting of two consecutive elements. Afterward, we ask the model for the first element of the list that contains a specific element. At this point, the specific element is the last element of a list. For example, when there are two lists with 3 elements each, the context and question are structured as follows:
@@
Since the elements of each list are presented in sequential order from the first to the last, the process of finding the list that contains the last element and then searching for the first element of that list must be performed in a right-to-left order. This poses a challenge for models that typically process context in a left-to-right order. Moreover, since this type of task does not involve specific factual knowledge, it minimizes the interference of the model’s pre-existing factual knowledge (Xiong et al., 2024). Thus, these characteristics of the synthetic task enable us to establish an environment that clearly exposes the model’s inherent misordered context problem. We create a synthetic dataset consisting of 1000 samples for the evaluation. Each sample is composed of 10 lists, each containing 3 integers between 0 and $9$ that do not overlap.
Models HotpotQA 2WikiMultihopQA MuSiQue
Llama-3.1-8B 49.31 61.19 38.10 55.77 40.75 34.86 32.05 26.24 27.06
+ CoRe ($\hat{k}=2$) 46.93 66.82 48.15 58.13 44.81 65.11 39.10 38.23 34.08
+ CoRe ($\hat{k}=3$) 50.46 70.54 50.74 62.58 47.44 68.18 43.64 41.02 36.08
Mistral-7B 34.42 41.14 18.82 32.90 29.77 15.43 15.14 10.31 11.00
+ CoRe ($\hat{k}=2$) 30.41 45.79 27.78 35.67 36.39 29.09 20.6 15.74 12.36
+ CoRe ($\hat{k}=3$) 42.50 54.51 38.93 42.91 38.67 28.70 23.96 20.17 14.78
Qwen2.5-7B 61.96 61.66 41.36 65.47 37.26 39.45 33.56 34.87 37.49
+ CoRe ($\hat{k}=2$) 67.03 70.33 54.20 70.30 42.40 64.54 44.76 43.10 45.08
+ CoRe ($\hat{k}=3$) 63.37 71.59 54.13 69.64 44.30 68.00 46.12 42.84 45.88
Phi-3.5-mini 26.44 44.17 27.70 45.26 27.00 35.19 17.60 16.80 12.67
+ CoRe ($\hat{k}=2$) 23.29 49.67 37.64 50.92 32.34 36.71 28.60 26.31 26.59
+ CoRe ($\hat{k}=3$) 41.41 59.07 40.62 46.91 33.76 22.98 31.49 26.83 32.40
GPT-4o-mini 66.92 68.66 51.96 76.63 43.84 60.96 45.26 42.04 40.04
+ CoRe ($\hat{k}=2$) 64.72 72.67 57.43 79.14 52.03 66.96 52.47 49.47 42.53
+ CoRe ($\hat{k}=3$) 66.52 74.06 56.72 79.84 51.09 68.84 52.66 49.37 42.89
We conduct experiments using a total of five models, including GPT-4o-mini (OpenAI, 2024), in addition to the four LLMs used in section 3.1. For Multi-hop QA tasks, we evaluate by measuring the F1 score based on the short answers generated by the models. We evaluate synthetic tasks by measuring accuracy based on whether the answers generated by the models match the correct answers. All prompts used in the experiments can be found in the Appendix C.
Figure 4: Main results in the synthetic task. Repetition step denotes the number of additional repetitions ($\hat{k}-1$).
@@ 4.2 Main Results
Table 1 shows the F1 scores of baseline models and models with CoRe applied for multi-hop QA tasks. Overall, performance improvements were observed across all models and datasets. Notably, the performance appears to reach a plateau in the $\hat{k}=2$ setting, where the context is repeated only once. Furthermore, we find additional performance improvements when further repetitions are applied ($\hat{k}=3$). In the case of HotpotQA, a significant performance improvement was seen in the bridge type. This is because answering bridge-type questions requires reasoning across documents by linking information between them, which is closely related to the order of documents. For 2WikiMultihopQA, a considerable performance improvement was observed in the bridge-comparison type, with Llama achieving a 30-point improvement and Qwen achieving a 25-point improvement in the F1 score. This is because the bridge-comparison type is a complex 4-hop reasoning task that combines bridge and comparison from HotpotQA. The MuSiQue dataset’s overall performance improved regardless of the number of hops. These results indicate that CoRe substantially enhances the model’s multi-hop reasoning performance, particularly in tasks where the order of supporting documents is crucial, effectively guiding the model along the correct reasoning paths.
@@
We measure the accuracy of each model as we increase the number of context repetitions up to 10 ($\hat{k}=11$) in the synthetic task. As shown in Figure 4, we can see that the accuracy when no repetition is performed has a low value between 0% and 60%. This demonstrates that existing LLMs are unable to perform the synthetic task directly at once. However, when CoRe is applied, we observe a significant performance improvement, with accuracy increasing by at least 30% points and up to 70% points. Additionally, we can see that many models tend to improve performance as the number of repetitions increases. This suggests that the model leads to the more appropriate order of reasoning chains as repetition progresses multiple times. However, for some models, particularly Qwen, there is a tendency for performance to drop again as the repetition steps increase. This appears to be due to noisy information being repeated in the context, negatively affecting its reasoning ability. We provide a detailed analysis of this in Section 4.3.